### v2

# Topological columnar nano-SQUID based on a 3D topological insulator

Abstract
Finding an ideal platform to generate non-Abelian Majorana zero-modes (MZMs) is a crucial first step for topological quantum computing. A 3D topological insulator (TI) is promising for this purpose due to the spin-momentum-locked surface state. Here we propose a novel yet simple TI platform that gives rise to a robust topological phase and can realize a well-defined fermion parity necessary for quantum computing. It consists of a bulk-insulating rectangular TI nanowire laterally sandwiched by two superconductors. In this structure, the top and bottom surfaces individually work as SNS line junctions, forming a nanometer-scale columnar SQUID in which the nanowire cross-section defines the threading flux $\Phi$ in axial magnetic fields. We demonstrate that a TI device of this structure indeed presents SQUID-type oscillations of the critical current $I_{c}$ as a function of $\Phi$ with period $\Phi_{0}^{s}=\frac{h}{2e}$. It can be tuned by a back gate to attain a full $I_{c}$ suppression and then a nearly-full recovery as a function of $\Phi$, an indication that the superconductivity is completely surface-dominated. Our theory shows that, when the two junctions are asymmetric, a robust topological phase hosting MZMs occurs periodically for $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$.
I Introduction
The main feature of three-dimensional (3D) topological insulators (TIs) is the existence of gapless surface Dirac fermions that are protected by time-reversal symmetry [1, 2, 3]. These Dirac fermions are spin-non-degenerate with the spin direction dictated by the momentum; this spin-momentum locking is the key ingredient for realizing “spinless” topological superconductivity that gives rise to Majorana zero-modes (MZMs) [4]. The non-Abelian statistics obeyed by MZMs allows for topological quantum computing based on braiding operations [5], so TIs have a high potential for such applications. However, TIs have been left behind in the main-stream efforts to realize topological quantum computing, largely due to the less developed technologies for TI device fabrications compared to semiconductor devices [6]. Unfortunately, despite significant efforts in the main-stream semiconductor platforms, the generation of MZMs is still elusive [7, 8]. This is mainly due to the sensitivity of the semiconductor platforms to Coulomb disorder [9, 8, 10, 11], and a new platform that is inherently insensitive to chemical-potential fluctuations is strongly called for.
In this regard, Josephson junctions (JJs) are an interesting platform. A JJ consisting of a normal section in-between two superconducting (SC) electrodes is called an SNS junction. The properties of such a JJ are governed by the Andreev bound states (ABSs) that form in the normal metal as a result of multiple Andreev reflections. Supercurrent flows through an SNS junction because Cooper pairs can be transported via the ABSs. In the original proposal by Fu and Kane [4] to realize MZMs in TIs, they considered an SNS line junction made on the TI surface. The ABSs in this junction become a 1D Majorana state that is $4\pi$-periodic as a function of the phase difference across the junction, $\varphi$, being topological for $\pi<\varphi<3\pi$.
Such a topological 1D Majorana state should in principle have a MZM when it has an end, but if the line junction ends in the middle of the TI surface, the MZM will hybridize with the gapless Dirac fermions in the rest of the surface; therefore, Fu and Kane proposed a phase-biased tri-junction to generate a MZM that is protected by a gap [4]. Even though a protected MZM can be generated in this way, when the tri-junction is made on the top surface of a TI and the bottom surface is left normal (which is expected when the TI thickness is more than $\sim$20 nm [12, 13]), the gapless Dirac fermions at the bottom surface may compromise the parity of the device, causing problems for a Majorana qubit that encodes quantum information through fermion parity [14]. Similar problem exists when using the MZM appearing at the Josephson vortex core of a TI line junction [15, 16]. Therefore, for topological quantum computing [5], gapping out the whole surface is important [17]. If this is achieved, the MZMs generated in a junction by phase biasing is robust and disorder-resilient [9], as their existence does not depend on the chemical potential at all provided that the TI is bulk-insulating.
Figure 1: Sandwich TINW Josephson junction. (a) False-color SEM image of device A; an exfoliated TI flake is dry-etched to leave a NW (pink), and the etched area is subsequently filled with Nb (green), such that the TINW is side-contacted by superconducting Nb. There are remaining TI pieces that do not play any role. (b) Schematic of our device: Back-gated TINW is laterally sandwiched between two superconducting Nb electrodes to a Josephson junction. In- and out-of-plane magnetic fields can be applied. (c) $dV/dI$ vs $I_{\rm b}$ data in device A at various $T$ shown for zero back-gate voltage; inset shows the $T$-dependence of the critical current $I_{c}$ extracted from these data. We define $I_{c}$ as the bias current above which $dI/dV$ exceeded 50 $\Omega$. (d) Differential resistance $dV/dI$ vs junction voltage $V$ for device A in $B$ = 0 T. Red arrows mark the position of the peaks due to multiple Andreev reflection (MAR). We assign the index $n$ from 1 to 4 to the peaks ($n$ = 1 for the largest $V$). Inset shows a plot of the $V$ value for index $n$ (which we call $V_{n}$) vs $1/n$. (e-g) Color mappings of $dV/dI$ as a function of $I_{\mathrm{b}}$ and $B_{\mathrm{z}}$ for devices A, B and C measured without any in-plane magnetic field at $T=30$ mK, presenting Fraunhofer patterns. The critical current manifests itself as a bright orange line to border the black area corresponding to zero resistance. Irregular Fraunhofer features are attributed to an inhomogenous supercurrent density along the junction.
To circumvent the problems and exploit the benefits of the TI line junction, we fabricate unconventional devices in which a rectangular nanowire, etch-fabricated from a bulk-insulating TI flake [18], is side-contacted by a superconductor on both left and right sides, creating a lateral sandwich SNS junction with a TI nanowire (TINW) as the N part (see Fig. 1b). In such a sandwich junction, if the TINW is bulk-insulating, the supercurrent flows only through the top and bottom surfaces (and a very small portion through the ends of the TINW) via the ABSs formed there. The whole surface of this TINW is gapped, with the proximity-induced SC gap in the side surfaces and the mini-gap of the ABSs elsewhere [4]. Since the axial magnetic field applied parallel to the TINW creates a magnetic flux $\Phi$ which controls the phase difference $\varphi$ at the top and bottom JJs, this structure can be viewed as a nanometer-scale columnar SQUID (superconducting quantum interference device) where the loop area corresponds to the TINW cross-section. A similar nano-SQUID was previously realized with a double-layer graphene heterostructure [19]. Our device has a bottom gate, which allows us to control the chemical potential in the TINW as well as to introduce asymmetry between the top and bottom junctions. A similar lateral sandwich junction was previously realized with a TI by using a diffusion process [20], but the device was not gate-tunable and it was not clear if the supercurrent flows only through the surface or also through the bulk.
In our devices, we find that the critical current $I_{c}$ of the device exhibits pronounced oscillations as a function of $\Phi$ with a period $h/(2e)$ and, in certain situations, $I_{c}$ is almost completely suppressed when $\Phi=h/(4e)$ but recovers almost entirely for $\Phi=h/(2e)$. This is exactly the behavior expected for a nano-SQUID with completely surface-dominated supercurrent with negligible bulk contribution. Theoretically, we first discuss a phenomenological SQUID model to explain the observed behavior, and then present numerical simulations based on a microscopic model of two line junctions made on a TINW surface, which support the SQUID model. Furthermore, we predict that when the SQUID is asymmetric, the magnetic flux $\Phi$ causes only one of the two line junctions to go through the topological phase transition of the Fu-Kane type [4] and the SQUID is topological for $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$. Note that there have also been related theoretical proposals to use the quasi-one-dimensional states of TINWs under superconducting proximity effect for generating MZMs [21, 22, 23, 24, 25]; however, these rely on the realization of a well-defined subband structure, i.e., that electron states on the top and bottom surfaces are coherently connected, which is most likely not realized in our device.
Figure 2: Behaviour as a nano-SQUID. (a-c) $dV/dI$ as a function of $I_{\mathrm{b}}$ and $B_{\mathrm{x}}$ obtained at $V_{\mathrm{G}}=0$ V in devices A, B and C. The $B_{z}$-component due to misalignment is compensated to realize a purely in-plane magnetic field. There can be $\sim$1 misalignment towards the $y$ axis, but a small $B_{y}$ component does not affect the data nor the interpretation. The sweep direction of $I_{\mathrm{b}}$ is different for positive and negative currents, such that the sweep always starts from zero. (d) $dV/dI$ as a function of $I_{\mathrm{b}}$ and $V_{\mathrm{G}}$ for device A. Coloured vertical lines mark the positions where the $I_{\mathrm{c}}(B_{x})$ behavior in panels (e) and (f) were measured. (e,f) $I_{\mathrm{c}}$ as a function of $B_{x}$ at various $V_{\mathrm{G}}$ values in devices A and B. All the data were acquired at $T=30$ mK.
II Results
Device characterization: Our TINWs are based on bulk-insulating BiSbTeSe2 [26, 27] and are side-contacted by Nb electrodes. A false-color scanning electron microscope (SEM) image of our device A is shown in Fig. 1a, while a device schematic is depicted in Fig. 1b along with the coordinate system. Note that the supercurrent flows across the TINW in our devices, while it flows along the wire in conventional devices [28, 29]. We observed robust supercurrents at low temperature (Fig. 1c) and the plot of differential resistance $dV/dI$ vs junction voltage $V$ (Fig. 1d) presents signatures of multiple Andreev reflection (MAR). The energy positions of these MAR peaks (Fig. 1d inset) give the SC gap energy $\Delta_{\rm MAR}\approx 0.9$ meV, which is close to the SC gap of the Nb electrodes (1.0 meV). From the $I$-$V$ characteristics, we can also extract [30, 31] the averaged junction transparency $\tau_{\rm av}$; our analysis (see supplement) gives $\tau_{\rm av}\approx$ 0.76, which is reasonably high for a TI-JJ [32]. These results demonstrate that the side-contact of a superconductor to a thin TI flake is possible and useful for the SC proximity effect. Similar side-contact technique is widely used for graphene [33, 34], but it is new to TI flakes.
Importantly, our sandwich TINW junction forms a well-defined Josephson junction along the full length. This is evidenced by the Fraunhofer patterns shown in Figs. 1e-g for three devices. Here, the magnetic field $B_{z}$ is applied perpendicular to the junction plane. The $B_{z}$-dependence of $I_{c}$ presents clear nodes, which strongly suggests that the supercurrent is not due to superconducting shorts but the SC phase difference $\varphi$ winds along the junction (in the $x$-direction) in response to $B_{z}$. In all three devices, the $B_{z}$ values at the first $I_{c}$ minimum are essentially consistent with the expected $B_{z}$ values to create the SC flux quantum $\Phi_{0}^{s}=h/2e$ in the junction area, taking into account the flux-focusing effect [32, 29, 35] (see supplement for details). We note that the observed Fraunhofer pattern is not very regular in all three devices. This is understood as a result of inhomogeneity in the critical current along the junction [32, 36, 37]. This irregularity suggests that either the side of the etched TINW is not smooth or the induced pairing potential is nonuniform, resulting in locally varying $I_{c}$.
Critical-current oscillations in parallel magnetic fields: To check for the SQUID behavior, we need to apply the magnetic field parallel to the TINW. This requires an accurate alignment, which was done by measuring $I_{c}$ as a function of small $B_{\mathrm{z}}$ in the presence of large nominal $B_{x}$ fields generated by our 6-1-1 T vector magnet (see supplement for details). By determining the necessary $B_{z}$ values to compensate for the misaligned $B_{z}$ component for large nominal $B_{x}$ values, we were able to completely cancel $B_{z}$ and obtain the $B_{x}$-dependence of the critical current in the absence of $B_{z}$.
Figures 2a-c show the $I_{c}(B_{x})$ behavior thus obtained in devices A, B, and C. One can see that $I_{c}$ presents pronounced oscillations as a function of $B_{x}$, almost vanishing at the minima in devices A and C, but virtually recovers its zero-field value at the maxima. These strong oscillations of $I_{c}$ are completely different from the Fraunhofer pattern and are akin to the SQUID pattern. To understand the period $\Delta B_{x}$ of these $I_{c}$ oscillations, we evaluated the geometrical and effective cross-sectional area of a total of five TINW devices: The geometrical area $A_{\mathrm{geo}}$ is calculated with the actual width $W$ and the thickness $d$, while the effective area $A_{\rm eff}$ takes into account the penetration depth of the surface-state wavefunction; namely, we assumed that the effective electronic boundary of the nanowire is $\sim$2.5 nm deep from the outer surface [38, 39, 40].
Figure 3: Relation between the TINW cross-section and the $B_{x}$-period of the $I_{c}$ oscillations. The period $\Delta B_{x}$, defined as the separation between the first minima for positive and negative $B_{x}$, is obtained from the $I_{c}$ oscillations in devices A-E (see supplement). The geometrical cross-section $A_{\mathrm{geo}}$ is simply $W$ times $d$. The effective cross-section $A_{\rm eff}$ takes into account the depth of the surface-state wavefunction ($\sim$2.5 nm) [38, 40]. $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$ with open and filled symbols, respectively. The solid and dashed lines are the expected relations when the period corresponds to the magnetic flux of $h/(2e)$ and $h/e$, respectively.
In Fig. 3, $A_{\mathrm{geo}}$ and $A_{\mathrm{eff}}$ are plotted vs $1/\Delta B_{x}$, along with the lines showing the area corresponding to the flux change $\Delta\Phi=h/e$ and $\Delta\Phi=h/(2e)$. One can easily see that the observed $\Delta B_{x}$ corresponds to the flux change of $h/(2e)$ if $A_{\rm eff}$ is taken. Thus, we conclude that the period of the $I_{c}$ oscillations is set by $h/(2e)$, which is expected for a SQUID. Note that the flux-focusing effect from the SC electrodes can be neglected in this geometry, because the Nb electrodes are almost fully penetrated by parallel magnetic fields. Note also that at the maximum $B_{x}$ field of this experiment, $\pm$6 T, the superconductivity of the Nb electrodes was about to be suppressed, making the $I_{c}$ lobes in Figs. 2a-c to be distorted towards 6 T.
It is useful to mention that a similar SQUID pattern in the $I_{c}$ oscillations due to topological edge states has been reported for a 2D TI, HgTe [41], and for a higher-order TI, Bi [42]; in these cases, a planar junction structure resulted in a SQUID behavior due to the dominance of supercurrents through the 1D edge states. In our case, the SQUID behavior owes to the dominance of 2D surface states in a sandwich junction structure. Note that the necessary magnetic field to generate $h/(2e)$ flux in the SQUID-loop is significantly larger in our device.
One of the advantages of our geometry is that the chemical potential $\mu$ can be tuned by a back-gate voltage $V_{\rm G}$. The $V_{\rm G}$-dependence of $I_{c}$ shown in Fig. 2d for device B demonstrates that $\mu$ can be tuned across the Dirac point, which appears to be located at around $V_{\rm G}\approx-10$ V. The $I_{c}$ increases by a factor of 3 as $\mu$ moves away from the Dirac point into the $n$-type regime, while $I_{c}$ presents a saturation in the $p$-type regime. Such a behaviour has been observed in planar JJs based on TIs [32, 43] and graphene [44], where the saturation in the $p$-type regime was attributed to the formation of a $pn$-junction between the gated junction channel and the areas that are in direct contact with the SC electrodes. The weak oscillatory $I_{c}(V_{\rm G})$ behavior in the $p$-type regime could be due to a Fabry-Perot-like interference [34, 44].
One can see in Figs. 2e-f that the pronounced $I_{c}$ oscillations as a function of $B_{x}$ are stable with respect to changes in $V_{\rm G}$, although an offset in $I_{c}$ grows with $V_{\rm G}$ in the $n$-type regime. Interestingly, device B is tuned with $V_{\rm G}\approx-20$ V to a regime where $I_{c}$ becomes zero at the minima. Such a disappearance of $I_{c}$ is possible only when there is no bulk contribution and the critical current at the top and bottom surfaces are equal. Hence, one can conclude that our TI nano-SQUID is surface-dominated.
Theoretical description: Let us start with a simple phenomenological model to describe the columnar TI nano-SQUID. We assume two independent JJs on the top and bottom of the TINW and further assume that the side surfaces are proximitized by Nb to have a finite SC gap $\Delta$ and they connect the two JJs. This is obviously a SQUID with the loop-area set by the cross-section of the nanowire. If we consider only the first harmonic in the current-phase relation (CPR) for each JJ, the total current of this SQUID is written as
$I(\varphi)\approx I_{t}\sin(\varphi-\pi\Phi/\Phi_{0}^{s})+I_{b}\sin(\varphi+\pi\Phi/\Phi_{0}^{s}),$ (1)
where $\varphi$ is the superconducting phase difference across the junctions and $I_{t}$ ($I_{b}$) is the maximum supercurrent through the top (bottom) surface. The critical current can be obtained by finding the maximum as a function of $\varphi$, $I_{c}=\max_{\varphi}\{I(\varphi)\}$. The resulting $I_{c}$ as a function of $\Phi$ are shown in Fig. 4a for two ratios of $I_{t}/I_{b}$. As is usual for a SQUID, $I_{c}$ exhibits large oscillations with periodicity $h/2e$, and the depth of the $I_{c}$ minima at $\Phi=h/4e$ depends on the asymmetry, $I_{t}/I_{b}$. We obtain $I_{c}\approx$ 0 at the minima in the symmetric case $I_{t}\approx I_{b}$. Therefore, the observed $V_{G}$ dependence of the $I_{c}$ oscillations shown in Figs. 2e-f is understood to be essentially a result of the gate-induced change in $I_{t}$ and $I_{b}$.
Figure 4: Results of numerical simulations: (a) The $I_{c}$ oscillations calculated for the simple phenomenological model (dotted lines) and the tight-binding model (solid line). The dominant period $\Delta\Phi=\Phi_{0}^{s}$ [$=h/(2e)$] appears in both models. In the phenomenological model Eq. (1), the ratio $I_{t}/I_{b}$ determines the size of the oscillations in $I_{c}$. The result of the tight-binding model for small $\Delta$ shows a superposition of oscillations with period $2\Phi_{0}^{s}$ [$=h/e$], which arises due to dependence of the normal state subbands on flux. This odd-even effect is not observed in our experiment, suggesting well-defined subbands is likely not formed. (b) Sketch: When the flux $\Phi$ is in the range $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ ($n$ odd), the columnar nano-SQUID hosts MZMs at the ends. (c) Topological phase diagram, spectral gap of top- and bottom surfaces $E_{{\rm gap,t/b}}$, and gauge-invariant equilibrium phase difference $\gamma_{0,{\rm t/b}}=\varphi_{0}-\frac{2\pi}{\Phi_{0}^{s}}\int{\bm{A}}{\rm d}{\bm{l}}=\varphi\pm\pi\Phi/\Phi_{0}^{s}$ at top- and bottom surfaces as function of magnetic flux $\Phi$ threading the nano-SQUID; calculated from Eq. (1). Topological regions are shaded in red.
To support this phenomenological nano-SQUID model, we performed numerical simulations of our set-up by using a BHZ tight-binding model of a 3D TINW with finite cross-section and of infinite length, such that the momentum $k$ along the wire is a good quantum number (see Methods). We assume a spatial dependence of the pairing potential $\Delta(y)$ induced in the TINW surface, such that there is a phase difference of $\varphi$ across the top/bottom surface; namely, we assume the SC phase of $-\varphi/2$ for the left side surface and $+\varphi/2$ for the right. The ground state value of $\varphi$ is set by the flux $\Phi$, but it changes when a transverse current flows. This model effectively describes a SQUID with two JJs at the top and bottom of the TINW surface. The original TINW states are now broken into four parts, two at the sides with a SC gap and two at the top and bottom with ABSs. By diagonalizing the Bogoliubov-de Gennes Hamiltonian for this tight-binding model, we calculate the total energy $E$ as a function of $\varphi$, and its derivative gives the CPR, which allows us to calculate $I_{c}$ (see Methods and supplement).
[…]
We note that the $h/e$ periodicity of the original TINW states is visible when $\Delta$ is small, but it disappears when $\Delta$ is large. This is because a large SC gap in the side surfaces effectively decouples the top and bottom surfaces (see supplement for details).
III Discussion
We now show theoretically that the nano-SQUID formed by the TI top and bottom surface states placed between the superconducting leads realizes a topological superconducting state over an extended flux range from $\Phi=(n-\frac{1}{2})\Phi_{0}^{s}$ to $(n+\frac{1}{2})\Phi_{0}^{s}$ around odd $n$ of flux quanta piercing the nanowire, see Fig. 4b for a sketch.
Recall that the spin-momentum locking in the TI surface states prohibits normal reflection at the NS interface for the perpendicular mode ($k$=0) and guarantees a pair of perfectly transmitted, non-degenerate ABSs (even- and odd-parity ABS) in the short-junction limit [4]. They present $4\pi$ periodicity and cross zero-energy at $\varphi=(2m+1)\pi$ with integer $m$, where the fermion parity of the ground state alternates [46, 47]. When this $(2m+1)\pi$ crossing occurs in only one of the junctions in a columnar nano-SQUID, it guarantees the appearance of a 1D topological phase [48, 4]. Now let us first consider the special case when an odd number of flux quanta pierces the nanowire, which leads to the superconducting phase to wind by an odd multiple of $2\pi$ in the nano-SQUID. The $2\pi$ relative phase difference (for example, phase zero at the top and $2\pi$ at the bottom) implies that the fermion parity on the top and bottom surfaces is opposite, making the SQUID topological. In this situation, the two ends of the columnar TI nano-SQUID become a boundary between the topological and trivial regions, which implies that MZMs should appear at the ends (Fig. 4b).
In a closed system, topological phase transitions occur every half-integer flux quanta, $\Phi=(n+\frac{1}{2})\Phi_{0}^{s},n\in\mathbb{Z}$. At these points, time-reversal symmetry (TRS) is restored when all effects of the magnetic field beyond the magnetic flux are neglected. Under this assumption, at $\Phi=(n+\frac{1}{2})\Phi_{0}^{s}$, due to TRS, one of the Josephson junctions must have a phase difference $\pi$, while the other has zero phase difference, which corresponds to the topological phase transition point in the SQUID. The asymmetry between top and bottom determines which surface turns topological: because a $\pi$ phase difference switches the sign of the Josephson energy, the junction with smaller Josephson energy acquires the $\pi$ phase difference. The topological region extends within a full flux quantum range $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ around each odd-integer $n$ of flux quanta threading the nano-SQUID, independent of the asymmetry between the top and bottom surfaces.
We verify the above picture by explicitly calculating the equilibrium phase difference from the nano-SQUID model Eq. (1) with asymmetry $I_{{\rm t}}/I_{{\rm b}}=0.8$. The result shown in Fig. 4c shows that the phase difference of the weaker top junction winds while the phase difference of the stronger bottom junction remains close to zero. Additionally, we show the spectral gap of top and bottom surfaces which is given by the energy of the perfectly transparent channel originating from the Dirac surface state, with functional form $E_{{\rm gap,t/b}}(\varphi)\propto\cos(\varphi/2)$ [4, 46]. The calculation confirms the spectral gap closure at the top surface at $\Phi=\Phi_{0}^{s}/2$ while the gap on the other surface remains open; this is the topological phase transition. Importantly, this calculation confirms the extended topological region for the flux bias, $(n-\frac{1}{2})\Phi_{0}^{s}<\Phi<(n+\frac{1}{2})\Phi_{0}^{s}$ with odd-integer $n$. In future, it is useful to perform more detailed analysis to understand the exact role of the front and back surfaces of the TINW as well as the exact location of the MZMs in the topological phase.
Our experiment demonstrates that the conceived columnar TI nano-SQUID can actually be fabricated with a new side-contacting technique and our device shows that the supercurrent is surface-dominated, which guarantees the prominent role of spin-momentum locking that dictates the appearance of perfectly-transmitted ABSs. Our theory shows that the occurrence of the topological phase transition in this nano-SQUID as a function of flux $\Phi$ is guaranteed by symmetry. Since the 1D topological superconductivity realized in a TI junction by phase-biasing is independent of chemical potential and the whole TI surface is gapped, our columnar nano-SQUID is a promising platform to host robust MZMs. To confirm the non-Abelian statistics of MZMs in this platform, it is important to realize a hard gap to completely get rid of low-energy quasiparticles and to theoretically develop concrete strategies to perform braiding.
Methods
Experimental details: The devices were fabricated on degenerately-doped Si substrates that were covered by a 290 nm SiO2 layer to enable the application of a global back-gate voltage. The TINWs are dry-etched with Ar plasma from mechanically-exfoliated thin flakes of the bulk-insulating TI material BiSbTeSe2 [26, 27]. They have a rectangular shape with width $W$ of 60 – 70 nm and a length $L$ of a couple of micrometers. We mainly report three devices with slightly varying dimensions; the full dimensions of the TINWs used for this work are summarized in the supplement. The TINWs are side-contacted by Nb electrodes ($T_{c}\sim 7$ K and $H_{\mathrm{c2,in\mbox{-}plane}}\gtrsim 6$ T) on two sides and along the full length of the wire. If not mentioned otherwise, all of the transport measurenents were performed at the base temperature ($\sim$30 mK) of our dry dilution refrigerator. We used a 6-1-1 T vector superconducting magnet for the precise alignment of the magnetic field along the TINW axis, which was crucial for this work. The differential resistance $dV/dI$ was measured with a lock-in technique at 40 Hz by superimposing a small AC current to the DC bias current $I_{\rm b}$.
Transport regime: Fitting the data in the inset of Fig. 1c to $I_{\mathrm{c}}\sim\sqrt{T}\exp(-2\pi k_{\mathrm{B}}T/E_{\mathrm{Th}})$ gives the Thouless energy $E_{\mathrm{Th}}\approx$ 1.5 meV. This is only slightly larger than the SC gap in Nb, implying that the junctions is in the intermediate to short junction regime [49]. The SC coherence length $\xi$ in the top or bottom surface is estimated as $\xi_{\mathrm{TI}}=\sqrt{\hbar D/\Delta}=\sqrt{E_{\mathrm{Th}}L^{2}/\Delta}\approx$ 75 nm ($L$ = 60 nm is the junction length). The $\xi$ on the side surface $\xi_{\rm side}$ should be much shorter, as the side surface is damaged by Ar etching. If we use $E_{\mathrm{Th}}=\hbar D/L^{2}$ with $D=v_{\mathrm{F}}l_{\mathrm{e}}/2$ for the estimate of the mean free path $l_{\mathrm{e}}$, we obtain $l_{\mathrm{e}}\approx 30$ nm for $v_{\mathrm{F}}=5.5\times 10^{5}$ m/s [27]. Hence, we find $l_{e}<L\lesssim\xi_{\rm TI}$.
Theory: To treat the TINW, we use a model for the 3D TI material BSTS [2, 23, 24, 50] with finite square cross-section of side length $12L+$ 1, such that the Bogoliubov-de Gennes Hamiltonian is given by
$H_{k}$ $=\frac{1}{2}\sum_{\begin{subarray}{c}n=-L\\ m=-L\end{subarray}}^{L,L}\mathbf{c}_{n,m,k}^{\dagger}\cdot\left\{M(k)\tau_{z}+A\sin(k)\tau_{x}\sigma_{x}-\mu\right\}\eta_{z}\mathbf{c}_{n,m,k}$ (2)
$+$ $\sum_{\begin{subarray}{c}n=-L\\ m=-L\end{subarray}}^{L-1,L}\left\{\mathbf{c}_{n+1,m,k}^{\dagger}\cdot\left\{-B\tau_{z}+\frac{iA}{2}\tau_{x}\sigma_{y}\right\}e^{i\eta_{z}\phi_{m}}\mathbf{c}_{n,m,k}\right\}\eta_{z}$
$+$ $\sum_{\begin{subarray}{c}n=-L\\ m=-L\end{subarray}}^{L,L-1}\left\{\mathbf{c}_{n,m+1,k}^{\dagger}\cdot\left\{-B\tau_{z}+\frac{iA}{2}\tau_{x}\sigma_{z}\right\}\mathbf{c}_{n,m,k}\right\}\eta_{z}$
$+$ $\frac{1}{2}\sum_{\begin{subarray}{c}n=-L\\ m=-L\end{subarray}}^{L,L}\mathbf{c}_{n,m,k}^{\dagger}\cdot\Delta(y)\left\{\cos[\varphi(y)]\eta_{x}+\sin[\varphi(y)]\eta_{y}\right\}\mathbf{c}_{n,m,k}$
$+$ H.c.,
where $M(k)=M_{0}-2B[\cos(k)-3]$ and the Pauli matrices $\tau_{i}$, $\sigma_{i}$, and $\eta_{i}$ act in orbital, spin, and particle-hole space, respectively. To model the nano-SQUID, we take the pairing potential to be finite on either side of the TINW, but zero for sites in the center
$\Delta(y)=\begin{cases}0\;\;\;{\rm if}\;n\leq N\\ \Delta\;\;\;{\rm otherwise}\end{cases}$ (3)
and the superconducting phase is given by
$\varphi(y)=\begin{cases}\varphi/2\;\;\;{\rm if}\;n\geq 0\\ -\varphi/2\;\;\;{\rm if}\;n<0.\end{cases}$ (4)
Finally, the Peierls term $\phi_{m}$ is set by the flux $\Phi$ such that,
$\phi_{m}=\frac{4m\pi\Phi}{(2L+1)^{2}}.$ (5)
For simplicity we choose parameters $A=1$, $M=-1.5$, $B=0.5$, $L=4$, and $\mu=0.5$. We choose the pairing potentials $\Delta=0.02$, which is smaller than the subband gap, and $\Delta=0.25$, which is larger than the subband gap (see Fig. 4). We set $N$=3 such that only the side sites in the cross-section have an induced pairing potential. It should also be noted that, due to the finite extent of the wavefunctions into the nanowire, the value of flux corresponding to the normal-state flux quantum $\Phi=\Phi_{0}^{n}=h/e$ is obtained from the normal-state band structure (see supplement).
From the tight-binding model in Eq. (2) we can obtain the energy dispersion $E_{i}(k)$, where $i$ is a band index, for a given value of flux $\Phi$ and SC phase difference $\varphi$. The full energy-phase relation is then obtained from the energy of the occupied states
[…]
which assumes zero temperature. This energy-phase relation is then used to obtain the CPR (shown in Fig. 4a) as its $\varphi$-derivative. Note that Eq. (6) utilises a momentum cutoff $\Lambda$, which is set such that the CPR has converged.
Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01). E.N. acknowledges support by the Studienstiftung des deutschen Volkes. The work at Massachusetts Institute of Technology was supported by a Simons Investigator Award from the Simons Foundation.
Data availability: The data that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.14331680 and Supplementary Information.